OpenStudy (anonymous):

If c is the measure of the hypotenuse of a right triangle, find each missing measure. If necessary, round to the nearest hundredth. a = 60, b = ?, c = 100 a. 116.62 b. 80.00 c. 160 d. 40

OpenStudy (anonymous):

@AQ99

OpenStudy (anonymous):

@MeowLover17

OpenStudy (anonymous):

The sides of a right triangle follow the Pythagorean theorem \[a^2+b^2=c^2\]

OpenStudy (anonymous):

60^2 + b^2 = 100^2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

d

OpenStudy (anonymous):

is that rigt

OpenStudy (anonymous):

no

OpenStudy (anonymous):

b² = 100² - 60² b² = 6400

OpenStudy (anonymous):

60*2 =120 100*2 =200

OpenStudy (anonymous):

80 ??

OpenStudy (anonymous):

The 2s are exponents, not multiplication. 100² means 100*100 and 60² means 60*60

OpenStudy (anonymous):

Yes it's 80

OpenStudy (anonymous):

did you just subtract 200 - 120 and said that was the answer?